OpenStudy (anonymous):

Someone please explain! A circle has its center at (1, 4) and a radius of 2 units. What is the equation of the circle? (x - 1)2 + (y - 4)2 = 4 (x + 2)2 + (y + 4)2 = 2 (x + 1)2 + (y - 4)2 = 4 (x - 1)2 + (y - 4)2 = 2

myininaya (myininaya):

\[(x-h)^2+(y-k)^2=r^2\] (h,k) center r radius

myininaya (myininaya):

simply plug into that to get the answer :)

OpenStudy (nenadmatematika):

(x-p)^2+(y-q)^2=r^2 is the equation of the circle where C(p,q) is the center and r is the radius...would you know to use these facts and make your equation?

OpenStudy (anonymous):

(x - 1)2 + (y - 4)2 = 2 is this right?

myininaya (myininaya):

r is 2 so what is r^2?

OpenStudy (anonymous):

(x - 1)2 + (y - 4)2 = 4 so it would this?

OpenStudy (anonymous):

would be*

OpenStudy (nenadmatematika):

yes, correct :D

OpenStudy (anonymous):

Thanks you both! (:

OpenStudy (nenadmatematika):

yw :)